# Dependent Dirichlet processes via thinning

### v1

@@ 1 Introduction
where $\delta_{\theta_{j}}$ indicates a Dirac mass at the location $\theta_{j}$, and the probabilities are defined as $\omega_{1}=v_{1}$, $\omega_{j}=v_{j}\prod_{h<j}(1-v_{h})$ for $j>1$. The DDP extends this formulation by defining a collection of random probability measures $p_{x}$, indexed by a covariate $x\in\mathcal{X}$, through modifications to the stick-breaking representation of the DP. Specifically, either the weights, the atoms, or both are replaced with stochastic processes indexed by $x$. This construction preserves the desirable property that the marginal distribution at each covariate value remains a DP. Moreover, when the covariate indexing the collection of random probability measures is categorical, say $g$ with levels $\{1,\dots,G\}$, case considered throughout this work, the DDP serves as a convenient prior modeling data structured into distinct yet related groups, e.g. patients undergoing different treatments. A notable instance of MacEachern’s DDP is the single-atom DDP. As the name suggests, this specification maintains a common set of atoms across all covariate values, introducing dependence solely through the weights. This formulation has been widely studied, particularly for modeling dependence over time (Griffin & Steel, 2011) and across spatial locations (Duan et al., 2007; Fuentes & Reich, 2013). A related approach was proposed by Griffin & Steel (2006), with the dependence between DPs introduced by modifying the ordering of the Beta-distributed random variables in their stick-breaking construction. Models for dependent probability measures based on the idea of sharing some or all the atoms extend beyond the DDP framework. The hierarchical Dirichlet process (HDP, Teh et al., 2006) models a collection of probability distributions as distinct DPs, introducing dependence and shared atoms by modeling the common base measure as a DP itself. The common atoms model (CAM, Denti et al., 2023) employs a nested DP structure with a shared set of atoms to jointly cluster groups defined by covariate values and individual observations. The Griffiths-Milne dependent Dirichlet process (GM-DDP, Lijoi et al., 2014) takes a different approach that explicitly distinguishes between covariate-specific and shared atoms, by defining a vector of dependent DPs as the superposition of covariate-specific processes and a shared one.
@@
We introduce the thinned dependent Dirichlet process (thinned-DDP), a novel nonparametric prior for modeling collections of dependent random probability measures $(p_{1},\ldots,p_{G})$. This prior is derived by thinning the atoms of a common DP. Instead of directly selecting the atoms $\theta_{j}$ that contribute to the definition of a covariate-specific random probability measure $p_{g}$, the approach we propose modifies the common weights $\omega_{j}$ in (1) by multiplying the common random variables $v_{j}$ by covariate-specific thinning variables $\ell_{j,g}\in\{0,1\}$, which determine whether the corresponding atoms $\theta_{j}$ appear in the definition of $p_{g}$. This approach retains the simplicity and tractability of the original stick-breaking construction, enabling analytical study of the properties of the thinned-DDP, and facilitating the adaptation of computational methods originally developed for a single DP to multi-sample settings. Unlike the single-atom DDP, the HDP, and the CAM, where all atoms are shared by all distributions, the thinned-DDP allows atoms to be either shared among subsets of measures or specific to individual ones. Moreover, the thinned-DDP offers a more flexible dependence structure compared to the GM-DDP. While both approaches account for the presence of shared and non-shared atoms, the thinned-DDP, unlike the GM-DDP, allows atoms to be shared by specific subsets of covariate-specific processes rather than by all of them. Finally, we observe that the idea of tilting the definition of a nonparametric process via thinning, that is by skipping some of its atoms, has found application in some recent contributions: Lau & Cripps (2022) introduced thinned completely random measures to model dependent hazard functions; the quasi-Bernoulli stick-breaking process of Zeng et al. (2023) uses a related idea to improve estimation of the number of clusters; Bi & Ji (2024) applied a thinning mechanism to the HDP to explicitly identify which atoms are specific to populations or common across them.
@@ 2 Dependent Dirichlet processes via thinning
where $s_{j}=\sum_{h=1}^{j-1}\ell_{h,1}\ell_{h,2}$ and $q_{j}=\sum_{h=1}^{j-1}\mathbb{I}_{\{\ell_{h,1}\neq\ell_{h,2}\}}$ are, respectively, the number of shared and distribution-specific atoms in the first $j-1$ components of $\ell_{1:2}$.
@@
By combining Proposition 2 with the asymptotic analysis of the number of distinct values in samples from a DP (Korwar & Hollander, 1973), we find that the growth of $\mathbb{E}[K(n_{1},n_{2};\ell_{1:2})]$, as $n_{1}=n_{2}=n\to\infty$, is bounded below by $\log(n)$ and above by $2\log(n)$, with the bounds being sharp.
@@ 4 Thinned-DDP mixture models
Although our proposed framework is general, in model (4) and throughout the remainder of the paper, we focus on the Bernoulli thinning specification introduced in Section 3.1, as it yields the most straightforward computational scheme. Given the constructive definition in (2), we find it convenient to adapt computational strategies, originally devised for the DP, relying on its stick-breaking representation, to the partially exchangeable setting considered here. Notable examples from the literature include the blocked Gibbs sampler (Ishwaran & James, 2001), the slice sampler and its efficient variants (Walker, 2007; Kalli et al., 2011; Ge et al., 2015), and the retrospective sampler (Papaspiliopoulos & Roberts, 2008). These methods, commonly referred to as conditional samplers, operate by explicitly updating the process components within the sampler and rely on either stochastic or deterministic truncation. We outline a blocked Gibbs sampler for the thinned-DDP mixture model with Bernoulli thinning in (4). The conjugate specification of kernel $\phi$ and base measure $P_{0}$ allows for the analytical derivation of full conditional distributions for all model parameters. The main steps of our algorithm mirror those of a standard blocked Gibbs sampler for a Dirichlet process, with two key distinctions: (i) our approach explicitly updates the thinning sequences $\ell_{1:G}$, and (ii) the update of the common stick-breaking variables $v_{j}$ accounts for observations across all samples $y_{g}$, with $g=1,\dots,G$. The complete algorithm is detailed in Section E of the Supplementary Material.
@@ 5 Simulation studies
We conduct a numerical investigation of the posterior properties of the thinned-DDP mixture model defined in Section 4 using simulated data consisting of $G$ samples. Specifically, we assess the accuracy of the estimated partitions of the observations and that of the posterior density estimates. The study is divided into two parts, presented in Sections 5.1 and 5.2. First, to illustrate how the thinned-DDP framework enables information sharing across groups while accounting for heterogeneity, we compare its results to two extreme modeling approaches: complete pooling and no pooling. In the complete-pooling approach, the data structure is ignored by combining all $G$ samples into a single one and fitting a common DP mixture, thereby maximizing information sharing. This formulation corresponds to setting $\pi_{g}=1$ for all groups in the thinned-DDP. On the other hand, the no-pooling approach adopts independent DP mixture models for the $G$ samples, thus preventing any information sharing across them. Second, to assess the performance of the thinned-DDP, we compare its results with two state-of-the-art modeling approaches: the CAM of Denti et al. (2023) and the GM-DDP of Lijoi et al. (2014). Like the thinned-DDP, both the CAM and the GM-DDP can be used as nonparametric models for grouped data while also inducing a partition of the entire data vector. Additionally, efficient algorithms for posterior inference are available for CAM and GM-DDP, implemented in the R packages SANple (D’Angelo & Denti, 2023) and BNPmix (Corradin et al., 2021), respectively. These features ensure a fair and meaningful comparison with our proposed approach.
We simulated grouped data $y_{1:G}$, where each sample $y_{g}$, with $g=1,\ldots,G$, follows a finite mixture of univariate Gaussians. Specifically, each $y_{g}$ was drawn from either a three-component mixture with means at $(-5,0,5)$ and probabilities $(0.5,0.25,0.25)$ or a two-component mixture with means at $(5,10)$ and probabilities $(0.4,0.6)$. The component variances were fixed at 0.6 across all mixtures. To evaluate how the number of groups and their sample sizes influence posterior inference, we considered multiple configurations. First, we generated data from $G=2$ populations with varying sample sizes: $(n_{1},n_{2})=\{(10,30),(20,60),(30,90),(40,120)\}$. Next, we extended the analysis to $G=10$ populations with sample sizes $(n_{g^{\prime}},n_{g^{\prime\prime}})=\{(10,30),(20,60),(30,90),(40,120)\}$ with $g^{\prime}=1,\dots,5$ and $g^{\prime\prime}=6,\dots,10$. This setup resulted in total sample sizes $n$ ranging from 40 to 800. In both scenarios, $G=2$ and $G=10$, half of the samples were drawn from the three-component mixture and the other half from the two-component mixture. Each experiment was repeated on 50 independent datasets. The algorithm settings and prior specifications are described in detail in Section F of the Supplementary Material.
@@ 5.1 Comparison with complete-pooling and no-pooling mixtures
We begin by evaluating the models’ ability to estimate the data partition. To do so, we use the realizations from the posterior distribution of the cluster allocation variables to construct group-specific posterior similarity matrices, where each entry $(i,i^{\prime})$ represents the posterior probability that observations $i$ and $i^{\prime}$, with $i,i^{\prime}\in\{1,\dots,n_{g}\}$, belong to the same cluster within group $g$. Next, we obtain point estimates of the partitions by minimizing the variation of information loss (Wade & Ghahramani, 2018), implemented via the salso R package (Dahl et al., 2023, 2022). To assess accuracy, we compute the Adjusted Rand Index (ARI; Rand, 1971; Hubert & Arabie, 1985) for each group, measuring the agreement between the estimated and true partitions. As a summary measure, we report the average ARI, calculated as the mean of the group-specific ARIs. Figure 3 represents the distribution of the ARI across 50 replications for all combinations of $G$ and $n$. The no-pooling model highlights a key limitation: when the sample size is small, the lack of information sharing between groups significantly impacts performance, leading to a lower ARI in the first configuration. In contrast, both the complete-pooling and thinned-DDP mixtures demonstrate strong performance in estimating the data partition in all settings.
@@ A.1 Proof of Proposition 1
We derive the correlation between pairs of distributions distributed as a thinned-DDP. Since the thinned-DDP belongs to the class of multivariate species sampling processes introduced by Franzolini et al. (2025), we can specialize their Proposition 2.7 to our setting. Specifically, Proposition 2.7 states that, if $(p_{1},p_{2})$ are distributed as a multivariate species sampling process, then, for every measurable set $A$ such that $10<P_{0}(A)<$ 1,
@@ A.2 Proof of Proposition 2
where ~$\tilde{\ell}_{1:2}$ is such that the two sequences ~$\tilde{\ell}_{1}$ and ~$\tilde{\ell}_{2}$ are identical. The same definition of ~$\tilde{\ell}_{1:2}$ gives rise to the following two almost sure identities:
@@
=(7)${{=}}\mathbb{E}[K_{0}^{(1:2)}(n_{1},n_{2};\tilde{\ell}_{1:2})+K_{0}(n_{1},n_{2};\tilde{\ell}_{1:2})]-\mathbb{E}[K_{0}(n_{1},n_{2};\tilde{\ell}_{1:2})]$
=(11)${{=}}\mathbb{E}[K(n_{1},n_{2};\tilde{\ell}_{1:2})]$
=$\stackrel{{\scriptstyle(a)}}{{=}}\sum_{i=1}^{n_{1}+n_{2}}\frac{\alpha}{\alpha+i-1},$
where $(a)$ holds since the specification of ~$\tilde{\ell}_{1:g}$ implies a framework of complete pooling with a single sample $y$ of size $n_{1}+n_{2}$ distributed according to a DP with concentration parameter $\alpha$.
@@ A.4 Proof of Proposition 3
=$\stackrel{{\scriptstyle(a)}}{{=}}w-\sum_{r=0}^{n_{1}}\binom{n_{1}}{r}(-1)^{r}\frac{\Gamma(1+r)\Gamma(\alpha)}{\Gamma(1+\alpha+r)}\sum_{j=1}^{w}\frac{\alpha^{j}}{(\alpha+r)^{j-1}}$
=$\stackrel{{\scriptstyle(b)}}{{=}}w-\sum_{r=1}^{n_{1}}(-1)^{r}\binom{n_{1}}{r}\frac{\Gamma(1+r)\Gamma(\alpha)}{\Gamma(1+\alpha+r)}\frac{\alpha}{r}\frac{(\alpha+r)^{w}-\alpha^{w}}{(\alpha+r)^{w-1}}$
@@
=$\stackrel{{\scriptstyle(c)}}{{=}}\sum_{m=0}^{n_{1}}\mathbb{E}_{v}\left[\mathbb{E}[K^{(w+1):\infty}\mid B=m,v]\right]\mathbb{E}_{v}\left[\Pr(B=m\mid v)\right],$ (14)

### v2

@@ 1 Introduction
where $\delta_{\theta_{j}}$ indicates a Dirac mass at the location $\theta_{j}$, and the probabilities are defined as $\omega_{1}=v_{1}$, $\omega_{j}=v_{j}\prod_{h<j}(1-v_{h})$ for $j>1$. The DDP extends this formulation by defining a collection of random probability measures $p_{x}$, indexed by a covariate $x\in\mathcal{X}$, through modifications to the stick-breaking representation of the DP. Specifically, either the weights, the atoms, or both are replaced with stochastic processes indexed by $x$. This construction preserves the desirable property that the marginal distribution at each covariate value remains a DP. Moreover, when the covariate indexing the collection of random probability measures is categorical, say $g$ with levels $\{1,\dots,G\}$, case considered throughout this work, the DDP serves as a convenient prior modeling data structured into distinct yet related groups, e.g. patients undergoing different treatments. A notable instance of MacEachern’s DDP is the single-atom DDP. As the name suggests, this specification maintains a common set of atoms across all covariate values, introducing dependence solely through the weights. This formulation has been widely studied, particularly for modeling dependence over time (Griffin and Steel, 2011) and across spatial locations (Duan et al., 2007; Fuentes and Reich, 2013). A related approach was proposed by Griffin and Steel (2006), with the dependence between DPs introduced by modifying the ordering of the Beta-distributed random variables in their stick-breaking construction. Models for dependent probability measures based on the idea of sharing some or all the atoms extend beyond the DDP framework. The hierarchical Dirichlet process (HDP, Teh et al., 2006) models a collection of probability distributions as distinct DPs, introducing dependence and shared atoms by modeling the common base measure as a DP itself. The common atoms model (CAM, Denti et al., 2023) employs a nested DP structure with a shared set of atoms to jointly cluster groups defined by covariate values and individual observations. The Griffiths-Milne dependent Dirichlet process (GM-DDP, Lijoi et al., 2014) takes a different approach that explicitly distinguishes between covariate-specific and shared atoms, by defining a vector of dependent DPs as the superposition of covariate-specific processes and a shared one.
@@
We introduce the thinned dependent Dirichlet process (thinned-DDP), a novel nonparametric prior for modeling collections of dependent random probability measures $(p_{1},\ldots,p_{G})$. This prior is derived by thinning the atoms of a common DP. Instead of directly selecting the atoms $\theta_{j}$ that contribute to the definition of a covariate-specific random probability measure $p_{g}$, the approach we propose modifies the common weights $\omega_{j}$ in (1) by multiplying the common random variables $v_{j}$ by covariate-specific thinning variables $\ell_{j,g}\in\{0,1\}$, which determine whether the corresponding atoms $\theta_{j}$ appear in the definition of $p_{g}$. This approach retains the simplicity and tractability of the original stick-breaking construction, enabling analytical study of the properties of the thinned-DDP, and facilitating the adaptation of computational methods originally developed for a single DP to multi-sample settings. Unlike the single-atom DDP, the HDP, and the CAM, where all atoms are shared by all distributions, the thinned-DDP allows atoms to be either shared among subsets of measures or specific to individual ones. Moreover, the thinned-DDP offers a more flexible dependence structure compared to the GM-DDP. While both approaches account for the presence of shared and non-shared atoms, the thinned-DDP, unlike the GM-DDP, allows atoms to be shared by specific subsets of covariate-specific processes rather than by all of them. Finally, we observe that the idea of tilting the definition of a nonparametric process via thinning, that is by skipping some of its atoms, has found application in some recent contributions: Lau and Cripps (2022) introduced thinned completely random measures to model dependent hazard functions; the quasi-Bernoulli stick-breaking process of Zeng et al. (2023) uses a related idea to improve estimation of the number of clusters; Bi and Ji (2024) applied a thinning mechanism to the HDP to explicitly identify which atoms are specific to populations or common across them.
@@ 2 Dependent Dirichlet processes via thinning
where $s_{j}=\sum_{h=1}^{j-1}\ell_{h,1}\ell_{h,2}$ and $q_{j}=\sum_{h=1}^{j-1}\mathbb{I}_{\{\ell_{h,1}\neq\ell_{h,2}\}}$ are, respectively, the number of shared and distribution-specific atoms in the first j1j-1 components of $\ell_{1:2}$.
@@
By combining Proposition 2 with the asymptotic analysis of the number of distinct values in samples from a DP (Korwar and Hollander, 1973), we find that the growth of $\mathbb{E}[K(n_{1},n_{2};\ell_{1:2})]$, as $n_{1}=n_{2}=n\to\infty$, is bounded below by $\log(n)$ and above by $2\log(n)$, with the bounds being sharp.
@@ 4 Thinned-DDP mixture models
Although our proposed framework is general, in model (4) and throughout the remainder of the paper, we focus on the Bernoulli thinning specification introduced in Section 3.1, as it yields the most straightforward computational scheme. Given the constructive definition in (2), we find it convenient to adapt computational strategies, originally devised for the DP, relying on its stick-breaking representation, to the partially exchangeable setting considered here. Notable examples from the literature include the blocked Gibbs sampler (Ishwaran and James, 2001), the slice sampler and its efficient variants (Walker, 2007; Kalli et al., 2011; Ge et al., 2015), and the retrospective sampler (Papaspiliopoulos and Roberts, 2008). These methods, commonly referred to as conditional samplers, operate by explicitly updating the process components within the sampler and rely on either stochastic or deterministic truncation. We outline a blocked Gibbs sampler for the thinned-DDP mixture model with Bernoulli thinning in (4). The conjugate specification of kernel $\phi$ and base measure $P_{0}$ allows for the analytical derivation of full conditional distributions for all model parameters. The main steps of our algorithm mirror those of a standard blocked Gibbs sampler for a Dirichlet process, with two key distinctions: (i) our approach explicitly updates the thinning sequences $\ell_{1:G}$, and (ii) the update of the common stick-breaking variables $v_{j}$ accounts for observations across all samples $y_{g}$, with $g=1,\dots,G$. The complete algorithm is detailed in Section E of the Supplementary Material.
@@ 5 Simulation studies
We conduct a numerical investigation of the posterior properties of the thinned-DDP mixture model defined in Section 4 using simulated data consisting of $G$ samples. Specifically, we assess the accuracy of the estimated partitions of the observations and that of the posterior density estimates. The study is divided into two parts, presented in Sections 5.1 and 5.2. First, to illustrate how the thinned-DDP framework enables information sharing across groups while accounting for heterogeneity, we compare its results to two extreme modeling approaches: complete pooling and no pooling. In the complete-pooling approach, the data structure is ignored by combining all $G$ samples into a single one and fitting a common DP mixture, thereby maximizing information sharing. This formulation corresponds to setting $\pi_{g}=1$ for all groups in the thinned-DDP. On the other hand, the no-pooling approach adopts independent DP mixture models for the $G$ samples, thus preventing any information sharing across them. Second, to assess the performance of the thinned-DDP, we compare its results with two state-of-the-art modeling approaches: the CAM of Denti et al. (2023) and the GM-DDP of Lijoi et al. (2014). Like the thinned-DDP, both the CAM and the GM-DDP can be used as nonparametric models for grouped data while also inducing a partition of the entire data vector. Additionally, efficient algorithms for posterior inference are available for CAM and GM-DDP, implemented in the R packages SANple (D’Angelo and Denti, 2023) and BNPmix (Corradin et al., 2021), respectively. These features ensure a fair and meaningful comparison with our proposed approach.
We simulated grouped data $y_{1:G}$, where each sample $y_{g}$, with $g=1,\ldots,G$, follows a finite mixture of univariate Gaussians. Specifically, each $y_{g}$ was drawn from either a three-component mixture with means at (5,0,5)(-5,0,5) and probabilities $(0.5,0.25,0.25)$ or a two-component mixture with means at $(5,10)$ and probabilities $(0.4,0.6)$. The component variances were fixed at 0.6 across all mixtures. To evaluate how the number of groups and their sample sizes influence posterior inference, we considered multiple configurations. First, we generated data from $G=2$ populations with varying sample sizes: $(n_{1},n_{2})=\{(10,30),(20,60),(30,90),(40,120)\}$. Next, we extended the analysis to $G=10$ populations with sample sizes $(n_{g^{\prime}},n_{g^{\prime\prime}})=\{(10,30),(20,60),(30,90),(40,120)\}$ with $g^{\prime}=1,\dots,5$ and $g^{\prime\prime}=6,\dots,10$. This setup resulted in total sample sizes $n$ ranging from 40 to 800. In both scenarios, $G=2$ and $G=10$, half of the samples were drawn from the three-component mixture and the other half from the two-component mixture. Each experiment was repeated on 50 independent datasets. The algorithm settings and prior specifications are described in detail in Section F of the Supplementary Material.
@@ 5.1 Comparison with complete-pooling and no-pooling mixtures
We begin by evaluating the models’ ability to estimate the data partition. To do so, we use the realizations from the posterior distribution of the cluster allocation variables to construct group-specific posterior similarity matrices, where each entry $(i,i^{\prime})$ represents the posterior probability that observations $i$ and $i^{\prime}$, with $i,i^{\prime}\in\{1,\dots,n_{g}\}$, belong to the same cluster within group $g$. Next, we obtain point estimates of the partitions by minimizing the variation of information loss (Wade and Ghahramani, 2018), implemented via the salso R package (Dahl et al., 2023, 2022). To assess accuracy, we compute the Adjusted Rand Index (ARI; Rand, 1971; Hubert and Arabie, 1985) for each group, measuring the agreement between the estimated and true partitions. As a summary measure, we report the average ARI, calculated as the mean of the group-specific ARIs. Figure 3 represents the distribution of the ARI across 50 replications for all combinations of $G$ and $n$. The no-pooling model highlights a key limitation: when the sample size is small, the lack of information sharing between groups significantly impacts performance, leading to a lower ARI in the first configuration. In contrast, both the complete-pooling and thinned-DDP mixtures demonstrate strong performance in estimating the data partition in all settings.
@@ 6 Real data analysis
Acknowledgments
Bernardo Nipoti and Andrea Ongaro are supported by the European Union Next Generation EU funds, component M4C2, investment 1.1., PRIN-PNRR 2022 (P2022H5WZ9).
Data Availability
The data that support the findings of this study are openly available in the R package BNPmix (Corradin et al., 2021) at http://doi.org/10.32614/CRAN.package.BNPmix.
@@ A.1 Proof of Proposition 1
We derive the correlation between pairs of distributions distributed as a thinned-DDP. Since the thinned-DDP belongs to the class of multivariate species sampling processes introduced by Franzolini et al. (2025), we can specialize their Proposition 2.7 to our setting. Specifically, Proposition 2.7 states that, if $(p_{1},p_{2})$ are distributed as a multivariate species sampling process, then, for every measurable set $A$ such that $0<P_{0}(A)<1$,
@@ A.2 Proof of Proposition 2
where $\tilde{\ell}_{1:2}$ is such that the two sequences $\tilde{\ell}_{1}$ and $\tilde{\ell}_{2}$ are identical. The same definition of $\tilde{\ell}_{1:2}$ gives rise to the following two almost sure identities:
@@
[…]
$\stackrel{{\scriptstyle(a)}}{{=}}\sum_{i=1}^{n_{1}+n_{2}}\frac{\alpha}{\alpha+i-1},$
where $(a)$ holds since the specification of $\tilde{\ell}_{1:g}$ implies a framework of complete pooling with a single sample $y$ of size $n_{1}+n_{2}$ distributed according to a DP with concentration parameter $\alpha$.
@@ A.4 Proof of Proposition 3
$\stackrel{{\scriptstyle(a)}}{{=}}w-\sum_{r=0}^{n_{1}}\binom{n_{1}}{r}(-1)^{r}\frac{\Gamma(1+r)\Gamma(\alpha)}{\Gamma(1+\alpha+r)}\sum_{j=1}^{w}\frac{\alpha^{j}}{(\alpha+r)^{j-1}}$
$\stackrel{{\scriptstyle(b)}}{{=}}w-\sum_{r=1}^{n_{1}}(-1)^{r}\binom{n_{1}}{r}\frac{\Gamma(1+r)\Gamma(\alpha)}{\Gamma(1+\alpha+r)}\frac{\alpha}{r}\frac{(\alpha+r)^{w}-\alpha^{w}}{(\alpha+r)^{w-1}}$
@@
$\stackrel{{\scriptstyle(c)}}{{=}}\sum_{m=0}^{n_{1}}\mathbb{E}_{v}\left[\mathbb{E}[K^{(w+1):\infty}\mid B=m,v]\right]\mathbb{E}_{v}\left[\Pr(B=m\mid v)\right],$ (14)